OpenStudy (anonymous):

solve the equation using the zero product property. (x-2)(x+7)=0

OpenStudy (kinggeorge):

The zero product property basically says that if you have two numbers \(a, b\) such that \(a\cdot b=0\), then either \(a=0\) or \(b=0\).

OpenStudy (campbell_st):

you need to solve the equations x - 2= 0 and x + 7 = 0 this will result in one factor being zero.... and\[0 \times a = 0\]

OpenStudy (kinggeorge):

Using this, you know that \((x-2)\) and \((x+7)\) are two numbers that when multiplied together equal to 0. Thus, either \(x-2=0\) or \(x+7=0\) can you solve each of these individually?